OpenStudy (anonymous):

If w=x^2+y-z+sin t & x+y=t , then how to find part.dervivative of w wrt y ? plz click to find more.....

OpenStudy (anonymous):

\[w = x ^{2} + y - z + \sin(t)\] \[x + y = t\] Sorry which derivative are you looking for?

OpenStudy (anonymous):

we have to find \[(dough w/dough y)_{x,z}\].partial derivative

OpenStudy (anonymous):

I'm honestly not sure what you mean by that notation.

OpenStudy (anonymous):

there's no notation here, consider dough = d (dw/dx)\[_{x,z}\]

OpenStudy (anonymous):

(\[(dw/dx)_{x,z}\] what does this x,z mean here?

OpenStudy (anonymous):

sry , it is \[(dw/dy)_{x,z}\]

OpenStudy (anonymous):

use \partial to get \[\partial\] if you like

OpenStudy (anonymous):

thanks eb... can u temme wat foes 'x,z' mean here?

OpenStudy (anonymous):

\[(\partial w/ \partial y)_{x,z}\] --------this is the actual notation

OpenStudy (anonymous):

Okay I think that means find the derivative of w with respect to y, and evaluate at x,z? Not exactly sure

OpenStudy (anonymous):

i wanted to know what this x,z means...i guess those are independent variables? is it?

OpenStudy (anonymous):

w=x2+y−z+sin(t) x+y=t w=x2+y−z+sin(x+y) assuming w=f(x,y,z) (dw/dy) =0+1+0+cos(x+y) =1+Cos(x+y)